# AsymFormer: Asymmetrical Cross-Modal Representation Learning for Mobile Platform Real-Time RGB-D Semantic Segmentation

### v5

@@ III-C Cross Modal Attention-Guided Feature Correlation Embedding
The $W_{1}$ and $W_{2}$ are embedded into $V_{1}$ and $V_{2}$ by dot product. The fused feature $Fused_{2}$ is the concatenation of $V_{1}$ and $V_{2}$:
$Fused_{2}=\text{Cat}[W_{1}\cdot V_{1},W_{2}\cdot V_{2}]\$ (13)
Finally, the $Fused_{2}$ is adapted into output channel $C_{2}$ and added with residual connection $Fused_{1}$.
$Fused_{2}=\text{ConvBN}_{1\times 1}(Fused_{2})+Fuesed_{1}$ (14)
TABLE I: Ablation experiment results for different multi-modal feature fusion method.
@@
IV EXPERIMENT RESULTS AND ANALYSIS
IV-A Implementation Details
To evaluate our Real-Time semantic segmentation network design, we conduct a series of experiments on two widely-used datasets: NYUv2[5] (795 training and 654 testing RGB-D images) and SUNRGBD[6] (5825 training and 5050 testing RGB-D images).
@@ IV-A Implementation Details
We conduct the model training and testing on different platforms. For the training, we use Nvidia A100-40G GPU. For the evaluation and inference speed testing, we use Nvidia RTX 3090 GPU, Ubuntu 20.04, CUDA 12.0 and Pytorch 2.0.1. We apply data augmentation to all datasets by randomly flipping (p=0.5), srandom scales [1.0,2.0], random crop 480$\times$640 and random HSV. We adopt Mix Transformer[17] and ConvNext[16] backbones pretrained on ImageNet-1k[24]. In the augmented setting, both backbones are pretrained on ADE20k[23]. For the CMA implementation, we set down-sample rate $f$=1 in CMA-1 and $f$=2 in CMA-2,3. The MLP-decoder in AsymFormer has the same structure as Segformer and an embedding dimension of 256. We choose AdamW optimizer with a weight decay of 0.01. The initial learning rate is $5e^{-5}$ and we use a poly learning rate schedule $(1-\frac{iter}{max_{i}ter})^{0.9r}$ with a warm-up of 10 epochs. We train with a batch size of 8 for NYUv2 (500 epochs) and SUNRGBD (200 epochs). We employ cross-entropy as the loss function and do not use any auxiliary loss during training process. The evaluation metric is mean Intersection over Union (mIoU).
IV-B Model Assessment
This paper proposes two different models based on the AsymFormer framework with different backbones. The AsymFormer uses ConvNext-T[16] for RGB representation learning and MiT-B0[17] for fused feature processing. This model has 33.0 million parameters and 36.0 GFLOPs computational cost, and it can achieve 65 FPS inference speed on RTX 3090. Particularly, with mixed precision quantization, AsymFormer achieves a real-time inference speed of 79 FPS on the RTX 3090 GPU. It also achieves a real-time inference speed of 29.4 FPS on the Tesla T4 GPU, which has a more limited computation resources (65 TFLOPs FP16).
IV-C Ablation Experiment
@@
We conduct a series of ablation experiments on NYUv2 dataset to evaluate the effectiveness of the LAFS and CMA module. We set two common feature fusion methods as our comparative baseline: 1. Cat: This method directly concatenates two features and then uses convolution layers to adjust the channel numbers. Essentially, it is a pixel-wise fusion without feature selection. 2. SE+MHSA: This method combines the popular SE attention [34] and MHSA attention [21] for feature fusion. Here, SE is used for feature selection in the channel direction, while MHSA is employed for further feature extraction on the fused features. Specifically, we use an MHSA with 2 heads to align with CMA.
@@
In our experiments, the Cat fusion method, used as a baseline, achieved a segmentation accuracy of 47.0 mIoU and an inference speed of 77.5 FPS. When using LAFS for feature selection alone, we achieved a performance improvement of 2.1% while sacrificing only 1.8 FPS of inference speed. This demonstrates that LAFS is a highly efficient feature selection method that provides performance gains without significantly impacting inference speed. In comparison to the other baseline, Cat+MHSA, which resulted in an 11.8 FPS reduction in inference speed, only a 2.9% improvement in segmentation accuracy was achieved. This further highlights the efficiency of LAFS.
Furthermore, we conducted experiments to investigate whether incorporating the mining of self-similarity within the two modalities after feature selection could further improve segmentation accuracy. When using CMA alone, we observed a 2.6% improvement in segmentation accuracy but encountered a significant decrease in inference speed of 10.1 FPS. Compared to LAFS, CMA showed a more noticeable reduction in inference speed. In comparison to SE+MHSA, CMA achieved similar accuracy, but it is important to note that no feature selection was performed in this case, indicating that relying solely on mining multi-modal self-similarity can achieve accuracy similar to the traditional feature fusion baseline SE+MHSA.
@@
Finally, we combined LAFS with CMA (LAFS+CMA). Since LAFS had minimal impact on inference speed and served a different purpose than CMA, the network’s inference speed only decreased by 2 FPS while achieving a significant improvement of 7.1% compared to the baseline Cat. At this point, the inference speed of LAFS+CMA was similar to SE+MHSA, but with a 2.1% performance improvement. This validates our experimental hypothesis: by re-modeling feature selection and mining cross-modal self-similarity, we can enhance the segmentation performance of the network without sacrificing inference speed compared to existing models. This demonstrates that we have indeed improved the efficiency of the network.
IV-D Comparison With State-of-The-Arts
IV-D1 NYUv2 Comparison Results
@@
TABLE II: Comparison Results on NYUv2. * denotes ADE20k pretrain. The inference speed is tested on RTX 3090 platform, (480 $\times$ 640) inputs.
Method Year Backbone Params/M Pixel Acc. (%) mIoU (%) Inf Speed/FPS Inf Speed (%)
PSCNet-T 2022 Res50 40 - 45.4 47 72.3%
PSCNet-L 2022 Res50 52 - 46.2 30 46.2%
ACNet 2019 Res50 118 - 48.3 28 43.1%
RDFNet 2017 Res152 450 76 49.1 15 23.1%
SA-Gate 2021 Res50 65 76.8 50.2 35 53.8%
ESANet 2022 Res34-Nbt1D 34 - 51.6 32 49.2%
MSFNet 2022 Res50 49 - 52.2 25 38.5%
FRNet 2022 Res34 86 77.6 53.6 39 60.0%
PGDENet 2022 Res34 101 78.1 53.7 32 49.2%
CMX-B2 2022 Segformer-B2 67 78.7 54.1 24 36.9%
AsymFormer 2023 B0+T 33 78.5 54.1 65 100.0%
AsymFormer (Multi-Scale) 2023 B0+T 33 79.3 55.3 - -
AsymFormer (FP16) 2023 B0+T 33 78.5 54.1 79 121.5%
IV-D2 SUNRGBD Comparison Results
Table III reports the performance of AsymFormer on the SUNRGBD dataset. AsymFormer achieves competitive accuracy with 49.1 % mIoU. Since the input image size is the same as NYUv2, all methods have the same inference speed as shown in Table II. However, the advantage of AsymFormer is not as significant as in NYUv2 experiment. For example, AsymFormer improves 3.9 mIoU over SA-Gate[27] in NYUv2 dataset (54.1% vs 50.2% mIoU), but decreases 0.3 mIoU in SUNRGBD dataset (49.1% vs 49.4% mIoU). A similar performance degradation can be observed in CMX-B2 result which also uses Transformer based backbone. We conjecture that this phenomenon may be caused by low quality depth images in SUNRGBD dataset. The aim of our research is not to construct a state-of-the-art method that has a marginal mIoU improvement over other methods, but to construct a method that has a better performance-speed balance and is more suitable for robot platform. Given that AsymFormer still has faster inference speed than other methods, we consider this performance acceptable for AsymFormer.
TABLE III: Comparison Results on SUNRGBD. # denotes multi-scale test.
@@ IV-D2 SUNRGBD Comparison Results
IV-E Visualization
1) LAFS Attention Map: As shown in Figure 2, to demonstrate that LAFS performs better than CBAM [18] in selecting features in the spatial dimension, we visualized the spatial attention weights of both methods. It can be observed that LAFS provides better coverage of informative regions in the image while maintaining consistency within objects and preserving the integrity of edges.
Figure 2: Difference between CBAM and LAFS’s spatial attention map.
2) Semantic Segmentation Results: The Figure 3 demonstrates the segmentation results of AsymFormer on the NYUv2 dataset. As observed, while maintaining a significantly faster inference speed compared to other methods, AsymFormer achieves comparable semantic segmentation accuracy to mainstream approaches.
@@ IV-E Visualization
V CONCLUSIONS
In this work, we proposed AsymFormer, which aims to construct a less-redundant real-time indoor scene understanding system. To enhance efficiency and reduce redundant parameters, we implemented the following improvement: 1. We employed an asymmetric backbone that compressed the parameters of the Depth feature extraction branch, thus reducing redundancy. 2. We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights. This resulted in a significant performance improvement while almost maintaining the inference speed. 3. Expanding on feature selection, we integrated modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters. The experiments demonstrated that the AsymFormer surpassed existing SOTA methods in terms of speed, demonstrating our method achieved a balance between accuracy and speed. Moving forward, we will continue to optimize the modules and address issues such as self-supervised pre-training of the model, aiming for further improvements.

### v7

@@ 3.3.4 Representation Learning in Multiple Subspaces
Then, $W_{1}$ and $W_{2}$ embed information into $V_{1}$ and $V_{2}$ through dot products. The fused feature $Fused_{2}$ is the concatenation of $V_{1}$ and $V_{2}$ in the channel dimension:
$Fused=\text{Cat}[W_{1}\cdot V_{1},W_{2}\cdot V_{2}]\$ (5)
Finally, CMA converts $Fused_{2}$ into output channels $C_{2}$ using a linear layer and connects it with the residual connection $Fused$, resulting in CMA’s output feature.
Table 1: Ablation experiment results for different multi-modal feature fusion method.
@@
4 EXPERIMENT RESULTS
Table 2: Comparison Results on NYUv2. The inference speed is tested on RTX 3090 platform, (480 $\times$ 640) inputs. MS denotes Multi-Scale inference strategy.
Method Year Backbone Params/M mIoU (%) Real-Time Speed/FPS Speed (%)
CMX-B2 [42] 2022 Segformer-B2 67 54.1 $\times$ 24 36.9%
Token-Fusion [33] 2022 Token-Fusion(S) - 54.2 $\times$ - -
CMX-B2 (MS) [42] 2022 Segformer-B2 67 54.4 $\times$ - -
Multi-MAE [1] 2022 Vit-B - 56.0 $\times$ - -
CMX-B4 (MS) [42] 2022 Segformer-B4 140 56.3 $\times$ - -
Omnivore [10] 2022 Swin-L - 56.8 $\times$ - -
CMX-B5 (MS) [42] 2022 Segformer-B5 181 56.9 $\times$ - -
SA-Gate [5] 2021 Res50 65 50.2 $\checkmark$ 35 53.8%
ESANet [26] 2022 Res34-Nbt1D 34 51.6 $\checkmark$ 32 49.2%
PSCNet-L [9] 2022 Res50 52 46.2 $\checkmark$ 30 46.2%
PSCNet-T [9] 2022 Res50 40 45.4 $\checkmark$ 47 72.3%
PGDENet [49] 2022 Res34 101 53.7 $\checkmark$ 32 49.2%
FRNet [50] 2022 Res34 86 53.6 $\checkmark$ 39 60.0%
DFormer-S [37] 2023 DFormer-S 18.7 53.6 $\checkmark$ 50 76.9%
AsymFormer 2024 B0+T 33 54.1 $\checkmark$ 65 100.0%
AsymFormer (FP16) 2024 B0+T 33 54.1 $\checkmark$ 79 121.5%
AsymFormer (MS) 2024 B0+T 33 55.3 $\times$ - -
4.1 Implementation Details
@@
To evaluate our Real-Time semantic segmentation network design, we conduct a series of experiments on two widely-used datasets NYUv2[28] (795 training and 654 testing RGB-D images) and SUNRGBD[30] (5825 training and 5050 testing RGB-D images). We conduct the model training and testing on different platforms. For the training, we use Nvidia A100-40G GPU. For the evaluation and inference speed testing, we use Nvidia RTX 3090 GPU, Ubuntu 20.04, CUDA 12.0 and Pytorch 2.0.1. We apply data augmentation to all datasets by randomly flipping (p=0.5), random scales between 1.0 and 2.0, random crop 480$\times$640 and random HSV. We adopt ConvNext-T [21] backbones pretrained on ImageNet-1k[6]. For the Mix Transformer-B0 [35], no pre-trained weight has been used. The MLP-decoder in AsymFormer has the same structure as Segformer and an embedding dimension of 256. We choose AdamW optimizer with a weight decay of 0.01. The initial learning rate is $5e^{-5}$ and we use a poly learning rate schedule $(1-\frac{iter}{max_{i}ter})^{0.9}$ with a warm-up of 10 epochs. We train with a batch size of 8 for NYUv2 (500 epochs) and SUNRGBD (200 epochs). We employ cross-entropy as the loss function and do not use any auxiliary loss during training process. The evaluation metric is mean Intersection over Union (mIoU).
4.2 Ablation Experiment
@@
We conduct a series of ablation experiments on NYUv2 dataset to evaluate the effectiveness of the LAFS and CMA module. We set two common feature fusion methods as our comparative baseline: 1. Cat: This method directly concatenates two features and then uses convolution layers to adjust the channel numbers. Essentially, it is a pixel-wise fusion without feature selection. 2. SE+MHSA: This method combines the popular SE attention [14] and MHSA attention [31] for feature fusion. Here, SE is used for feature selection in the channel direction, while MHSA is employed for further feature extraction on the fused features.
@@
In our experiments, the Cat fusion method, used as a baseline, achieved a segmentation accuracy of 47.0 mIoU and an inference speed of 77.5 FPS. When using LAFS alone, we achieved a performance improvement of 2.1% while sacrificing only 1.8 FPS of inference speed. This demonstrates that LAFS provides performance gains without significantly impacting inference speed. In comparison to the other baseline, using Cat+MHSA, which resulted in a reduction of inference speed by 11.8 FPS, an improvement of only 2.9% in mIoU was achieved. This further highlights the efficiency of LAFS. Furthermore, when using CMA alone, we observed a 2.6% improvement in segmentation accuracy but encountered a significant decrease in inference speed of 10.1 FPS. Compared to LAFS, CMA showed a more noticeable reduction in inference speed.
@@
Finally, we combined LAFS with CMA (LAFS+CMA). Since LAFS had minimal impact on inference speed and served a different purpose than CMA, the network’s inference speed decreased by only 2 FPS. This change achieved a significant improvement of 7.1% in segmentation accuracy compared to the baseline Cat. At this point, the inference speed of LAFS+CMA was similar to SE+MHSA, but with a 4.2% performance improvement. This validates our experimental hypothesis: by re-modeling feature selection and mining cross-modal self-similarity, we can enhance the segmentation performance of the network without sacrificing inference speed compared to existing models. This demonstrates that we have indeed improved the efficiency of the network.
4.3 Comparison With State-of-The-Arts
4.3.1 NYUv2 Comparison Results
@@
4.3.2 SUNRGBD Comparison Results
Table 3 reports the performance of AsymFormer on the SUNRGBD dataset. AsymFormer achieves competitive accuracy with 49.1 % mIoU. The advantage of AsymFormer is not as significant as in NYUv2 experiment. For example, AsymFormer improves 3.9 mIoU over SA-Gate[5] in NYUv2 dataset (54.1% vs 50.2% mIoU), but decreases 0.3 mIoU in SUNRGBD dataset (49.1% vs 49.4% mIoU). A similar performance degradation can be observed in CMX-B2 result which also uses Transformer based backbone. We conjecture that this phenomenon may be caused by low quality depth images in SUNRGBD dataset. The aim of our research is not to construct a state-of-the-art method that has a marginal mIoU improvement over other methods, but to construct a method that has a better performance-speed balance and is more suitable for robot platform. Given that AsymFormer still has faster inference speed than other methods, we consider this performance acceptable for AsymFormer.
Table 3: Comparison Results on SUNRGBD. MS denotes Multi-Scale inference strategy.
@@ 4.3.2 SUNRGBD Comparison Results
4.4 Visualization
4.4.1 LAFS Attention Map
As shown in Figure 6, to demonstrate that LAFS performs better than CBAM [34] in selecting features in the spatial dimension, we visualized the spatial attention weights of both methods. It can be observed that LAFS provides better coverage of informative regions in the image while maintaining consistency within objects and preserving the integrity of edges.
Figure 6: Difference between CBAM and LAFS.
4.4.2 Semantic Segmentation Results
The Figure 7 demonstrates the segmentation results of AsymFormer on the NYUv2 dataset. As observed, while maintaining a significantly faster inference speed compared to other methods, AsymFormer achieves comparable semantic segmentation accuracy to mainstream approaches.
@@ 4.4.2 Semantic Segmentation Results
5 CONCLUSIONS
In this work, we proposed AsymFormer, which aims to construct a less-redundant real-time indoor scene understanding network. To enhance efficiency and reduce redundant parameters, we implemented the following improvement: 1. the asymmetric backbone that compressed the parameters of the Depth feature extraction branch, thus reducing redundancy. 2. the LAFS module for feature selection, utilizing learnable strategy for global information compressing and improving spatial attention calculations 3. the self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters. The experiments demonstrated that the AsymFormer achieves a balance between accuracy and speed. Moving forward, we will continue to optimize the modules and address issues such as self-supervised pre-training of the model, aiming for further improvements.